# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v3

@@ Abstract
Despite the availability of commercial QRNG devices, distinguishing between PRNG and QRNG outputs computationally remains challenging. This paper presents two significant contributions from the perspectives of QRNG manufacturers and users. For manufacturers, the conventional method of assessing the quantumness of single-photon-based QRNGs through mean and variance comparisons of photon counts is statistically unreliable due to finite sample sizes. Given the sub-Poissonian statistics of single photons, confirming the underlying distribution is crucial for validating a QRNG’s quantumness. We propose a more efficient two-fold statistical approach to ensure the quantumness of optical sources with the desired confidence level. Additionally, we demonstrate that the output of QRNGs from exponential and uniform distributions exhibit similarity under device noise, deriving corresponding photon statistics and conditions for $\epsilon$-randomness.
@@
From the user’s perspective, the fundamental parameters of a QRNG are quantumness (security), efficiency (randomness and random number generation rate), and cost. Our analysis reveals that these parameters depend on three factors, expected photon count per unit time, external reference cycle duration, and detection efficiency. A lower expected photon count enhances security but increases cost and decreases the generation rate. A shorter external reference cycle boosts security but must exceed a minimum threshold to minimize timing errors, with minor impacts on cost and rate. Lower detection efficiency enhances security and lowers cost but reduces the generation rate. Finally, to validate our results, we perform statistical tests like NIST, Dieharder, AIS-31, ENT etc. over the data simulated with different values of the above parameters. Our findings can empower manufacturers to customize QRNGs to meet user needs effectively.
Photon Statistics, Randomness, Quantumness, QRNG
I Introduction
@@
Random numbers have important roles in numerous applications, including computational methods like Monte Carlo simulations and programming Carlson et al. (2015); Metropolis and Ulam (1949), testing principles of physics and chemistry Bell (1964), as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett et al. (1992); Shannon (1949), extending to commercial uses such as slot machines and lottery games Konstantinou et al. (2004); Yamanaka and Nakano (2017). In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the contrary, the randomness of a true random number generator (TRNG) is rooted in some physical entropy source Schindler (2009); Stipčević and Koç (2014). A quantum random number generator (QRNG) uses the quantum particles and the principles of quantum physics Ma et al. (2016) to produce random numbers.
@@
Statistically, it is not feasible to distinguish between a PRNG and a QRNG just by analyzing their outputs. Statistical tests like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), AIS-$31$ Killmann and Schindler (2011), ENT Walker (2008) etc. are available to test the randomness of a random number generator (RNG). However, from the results of these tests, it is not possible to say whether an RNG is quantum or not. So a pertinent question arises:
@@
Different technologies, including photon arrival time Stipčević and Rogina (2007); Dynes et al. (2008); Wayne and Kwiat (2010); Michael A. Wayne and Kwiat (2009); Wahl et al. (2011); Nie et al. (2014); Khanmohammadi et al. (2015); Zhang et al. (2021), tunneling Bernardo-Gavito et al. (2017); Aungskunsiri et al. (2021), fluctuation Gabriel et al. (2010); Haw et al. (2015); Guo et al. (2019); Zhou et al. (2019); Zheng et al. (2019); Gehring et al. (2021), phase noise Qi et al. (2010); Guo et al. (2010); Shen et al. (2010); Xu et al. (2012); Nie et al. (2015); Yang et al. (2016); Huang et al. (2020), etc. have been proposed to produce QRNGs Herrero-Collantes and Garcia-Escartin (2017); Mannalatha et al. (2023). Several manufacturers such as ID Quantique ID Quantique (2001), Qutools Qutools (2010), QNuLabs QNu Labs (2018), QuintessenceLabs Quintessence Labs (2022), Mars Innovation Mars Innovation and Department of Electrical and Computer Engineering & Department of Physics, University of Toronto (2023), Quantum Computing Inc. Quntum Computing Inc. (2023), Quside Quside (2023), QuantumCTek Quantum CTek (2023) are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of the companies follow this approach, in this work we try to answer the Q1 with respect to this type of QRNGs.
In most of the analyses Jones et al. (2023); Peřina et al. (2024a, b); Bush and Børkje (2024) of the photon arrival-time-based QRNGs, the photon statistics have been discussed either by comparing the mean and variance of the sample distribution or using the negativity of the Mandel parameter Mandel (1979). For example, Alléaume et al. Alléaume et al. (2004) discussed photon statistics of photon sources using the Mandel parameter given by (Mandel, 1979, Eq. (11a)). They calculated values of the Mandel parameter from observed photon count timestamps and concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources. Although analytically it is true, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same as the mean and the variance of the underlying distribution. Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
Another approach to test the quantumness of light is to experimentally observe photon antibunching Paul (1982). However, although a stream of antibunched photons indicates the non-classical nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons Zou and Mandel (1990); Fox (2006). We explain this in more detail in Appendix A. This approach also consists of an equality criteria similar to above photon statistics case. Therefore this is also not enough to decide photon antibunching.
On the other hand, a single-photon source ensures the quantumness of a QRNG Paul (1982); Fox (2006). However, generating single photons are experimentally challenging taskImamoḡlu and Yamamoto (1994); Kim et al. (1999). Various experimental methods Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been explored to produce single photons. Existing time-bin-based QRNG models utilize coherent photon sources in conjunction with attenuators to produce random numbers from exponential Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Wahl et al. (2011) or uniform Stipčević and Rogina (2007) distributions if the underlying optical devices perform ideally and the source generates perfectly single photon. But, in practice, this is not the case. Therefore, it is important to explore the photon statistics in practical scenarios.
@@
Also, all of the QRNGs require some classical post-processing using deterministic functions. These post-processings use a classical source of randomness to increase randomness in the raw data generated by the QRNGs. Side-channel attacks Kuznetsov et al. (2019); Park et al. (2019); Chowdhury et al. (2021) can also be performed in this post-processing step to gather information about the random number. Therefore, the amount of classical post-processing should be minimal to ensure the randomness from the quantum source in the final outputs and reduce the possibility of side-channel attacks on classical post-processing. To achieve this, the raw output number (before classical post-processing) generated by QRNG should be as close to uniform random as possible. As all the practical devices are prone to error, this raises an important question:
Q2: How to increase the randomness from the quantum source in the presence of device errors?
@@
Besides quantumness and randomness, the random number generation rate and the cost of the QRNG are also important parameters from the user’s perspective. The trade-off between the randomness from the quantum source, the random number generation rate and the cost of the QRNG depends on the users’ requirements. For example, in the case of secure communication, the randomness from the quantum source should have the highest priority, and also the generation rate should be high. On the other hand, in token generation or lottery games, a low rate of random number generation would be sufficient. Therefore, to fulfill the user’s requirement, the answers to the following questions are required.
@@
In this article, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level.
@@
As another contribution, also from the manufacturer’s perspective, we derive the photon statistics for two-photon arrival-time-based QRNGs in practical scenarios. These photon statistics for QRNGs can be considered as a measure of randomness from quantum sources for photon arrival-time-based QRNGs. Depending on this measure, a suitable post-processing algorithm can be chosen to retain the maximum amount of randomness from the quantum source. For the practical scenario, we consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
@@
From the user’s perspective, we see that the maximum randomness from the quantum source may not be achieved due to some practical constraints related to the random number generation rate, the timing error in registering the photon detection time and the cost of the QRNG. In particular, we show the following.
@@
  • If the expected number of photons within a unit of time is small, the randomness from the quantum source would be high. However, it would increase the cost of the QRNG, and also, it would reduce the random number generation rate.
@@
  • Although a small external reference cycle of the QRNG increases the randomness from the quantum source, to minimize the timing error, it should be higher than a specific value. This reference cycle does not have much effect on the cost or random number generation rate of the QRNG.
@@
  • A detector with less efficiency increases the randomness from the quantum source and also reduces the cost of the QRNG. However, it also reduces the random number generation rate.
@@
Finally, we simulated random numbers depending on photon statistics we have derived, and performed statistical tests like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), AIS-$31$ Killmann and Schindler (2011) and ENT Walker (2008) on those random numbers. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources.
@@
This paper is structured as follows. In Section II we perform some preliminary discussion about photon statistics, some statistical tools used in this work and common randomness testing suites. In Section III and IV we revisit single-photon-based QRNGs and coherent pulse-based QRNGs respectively and mention their weakness in present days. Section V reveals our two-fold proposal to check the quantumness of a quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section VI to answer Q2. Section VII answers Q3 by separately analyzing the effects of the parameters mentioned in that question. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section VI are given in Section VIII. Finally, in Section IX we conclude our work.
II Preliminaries
In this section we first discuss photon statistics. We also discuss how these photon statistics distinguish quantum light from classical light. After that, we mention some statistical tools used in this work. Then we briefly discuss common randomness testing suites for testing the randomness of random numbers
II.1 Photon Statistics
@@
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols Bennett et al. (1992); Shannon (1949). In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, the photon-counting statistics have an important role in determining whether the light falls in the quantum regime or not.
@@
The distribution of the photon-counting statistics can be divided into three classes Fox (2006):
@@
Classical wave theory of light follows Poissonian and super-Poissonian statistics Fox (2006). On the other hand, the sub-Poissonian nature of a photon source is a clear indication of its quantumness Davidovich (1996); Paraïso et al. (2021).
@@
$P(n)=\begin{cases}\frac{e^{-\alpha\mu}(\alpha\mu)^{\alpha n}}{(\alpha n)!},\;n=0,1,2,\dots\,&\text{if }\alpha n\text{ is integer,}\\ 0,&\text{otherwise}\end{cases}$ (1)
for some positive real numbers $\mu,\alpha$. Then the mean would be $\mu$ and variance would be $\frac{\mu}{\alpha}$. Therefore, this denotes a sub-Poissonian, Poissonian, or super-Poissonian distribution for $\alpha>1,\alpha=1,\alpha<1$ respectively.
Thermal light follows Bose–Einstein distribution Fox (2006), given by
@@
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is the creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution Titulaer and Glauber (1966) given by
@@
II.2 Statistical Tools
II.2.1 Interval Estimation
II.2.2 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution. this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample Casella and Berger (2002). A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
A hypothesis test starts with two hypotheses: the null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{a}$). Depending on the significance level $\alpha$, the critical region is decided. The acceptance or rejection decision of $H_{0}$ is based on what is called the $p$-value of the test statistic. It is defined as $P(|X|>|x|)$, where $X$ is a random variable, whose distribution is the reference distribution of the test statistic of a hypothesis test, and $x$ is the observed value of this statistic Ross (2009); Hogg and Craig (2019). When the observed test statistic falls within the critical region, $p$-value of the test statistic becomes less than $\alpha$. In this case, $H_{0}$ is rejected. If the $p$-value of the test statistic is greater than $\alpha$, the observed test statistic falls outside the critical region. In that case, $H_{0}$ will be accepted.
II.2.3 Goodness-of-fit Test
$H_{0}$: Sample data comes from a distribution to be tested.
$H_{a}$: Sample data not from that distribution.
Chi-square ($\chi^{2}$) testCochran (1952), Kolmogorov–Smirnov test (KS test) Massey (1951), and Anderson–Darling test Anderson and Darling (1952) are some examples of the goodness-of-fit test Dodge (2008a). Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well Conover (1972). The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions Slakter (1965).
II.3 Randomness Testing Suites
Definition 1.
$X=x_{1}x_{2}\dots x_{n},x_{i}\in\{0,1\}\ \forall i,\text{ is true random}$
where $x=x_{j}x_{j+1}\dots x_{j+k}$.
To overcome this limitation, several randomness testing suites like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), AIS-$31$ Killmann and Schindler (2011), ENT Walker (2008) etc. have been proposed. Recently, Foreman et al. also suggested some test suite Foreman et al. (2024) for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation Metropolis and Ulam (1949) to validate the randomness.
III Revisiting QRNG based on Single-photon Source
QRNG, based on single photons, uses $/50$ 50 beam splitter Jennewein et al. (2000), where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Fig. 1.

### v4

@@ Abstract
This work presents two significant contributions from the perspectives of QRNG manufacturers and users. For manufacturers, the conventional method of assessing the quantumness of single-photon-based QRNGs through mean and variance comparisons of photon counts is statistically unreliable due to finite sample sizes. Given the sub-Poissonian statistics of single photons, confirming the underlying distribution is crucial for validating a QRNG’s quantumness. We propose a more efficient two-fold statistical approach to ensure the quantumness of optical sources with the desired confidence level. Additionally, we demonstrate that the output of QRNGs from exponential and uniform distributions exhibit similarity under device noise, deriving corresponding photon statistics and conditions for $\epsilon$-randomness.
@@
From the user’s perspective, the fundamental parameters of a QRNG are quantumness, efficiency (random entropy and random number generation rate), and cost. Our analysis reveals that these parameters depend on three factors, namely, expected photon count per unit time, external reference cycle duration, and detection efficiency. A lower expected photon count enhances entropy but increases cost and decreases the generation rate. A shorter external reference cycle boosts entropy but must exceed a minimum threshold to minimize timing errors, with minor impacts on cost and rate. Lower detection efficiency enhances entropy and lowers cost but reduces the generation rate. Finally, to validate our results, we perform statistical tests like NIST, Dieharder, AIS-31, ENT etc. over the data simulated with different values of the above parameters. Our findings can empower manufacturers to customize QRNGs to meet user needs effectively.
1 Introduction
@@
Random numbers have important roles in numerous applications, including computational methods like Monte Carlo simulations and programming [1, 2], testing principles of physics and chemistry [3], as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages [4, 5], extending to commercial uses such as slot machines and lottery games [6, 7]. In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the contrary, the randomness of a true random number generator (TRNG) is rooted in some physical entropy source [8, 9], which may be classical or quantum.
The randomness of a classical true random number generator (CTRNG) depends on the unpredictability of the physical process that generates the entropy. This unpredictability arises from our limited knowledge of the system’s behavior [9, 10]. However, if future research resolves these limitations, the process could become predictable and would no longer function as a true random number generator. In contrast, a quantum true random number generator (QTRNG, or more commonly, QRNG) exploits the properties of quantum particles and the principles of quantum mechanics [11] to produce random numbers. According to quantum theory, these numbers are fundamentally random, making them inherently unpredictable. Thus it would remain unpredictable, irrespective of any technological advancement.
Different technologies, including photon arrival time [12, 13, 14, 15, 16, 17, 18, 19], tunneling [20, 21], fluctuation [22, 23, 24, 25, 26, 27], phase noise [28, 29, 30, 31, 32, 33, 34], etc. have been proposed to produce QRNGs [35, 10]. Several manufacturers such as ID Quantique [36], Qutools [37], QNuLabs [38], QuintessenceLabs [39], Mars Innovation [40], Quantum Computing Inc. [41], Quside [42], QuantumCTek [43] are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of companies follow this approach, in this work we focus on the analysis of this type of QRNGs only.
Since no black-box randomness-testing algorithm can distinguish between a PRNG and a QRNG, therefore the first step in validating the quantumness [^1] Sub-Poissonian distribution ($mean>variance$) of photon statistics (more details in Section 2). of a random number generator (RNG) is to collect information regarding intermediate attenuator and detector parameters and photon statistics. Then it may be possible to devise a strategy to decide whether the device performs in the classical or quantum regime.
In most of the analyses [44, 45, 46, 47] of the photon arrival-time-based QRNGs, the photon statistics have been discussed either by comparing the mean and variance of the sample distribution or using the negativity of the Mandel parameter [48]. For example, Alléaume et al. [49] discussed photon statistics of photon sources using the Mandel parameter given by [48, Eq. (11a)]. They calculated values of the Mandel parameter from observed photon count timestamps and concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources.
Although the above strategy is analytically correct, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same (even under approximation) as the mean and the variance of the underlying distribution [50]. Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
Another approach to test the quantumness of light is to observe photon antibunching [51] experimentally. However, although a stream of antibunched photons indicates the non-classical nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons [52, 53]. We explain this in more detail in Appendix B. This approach also consists of equality criteria similar to the above photon statistics case. Therefore this is also not enough to decide photon antibunching.
On the other hand, a single-photon source ensures the quantumness of a QRNG [51, 53]. However, generating single photons are experimentally challenging task [54, 55]. Various experimental methods [56, 57, 58, 59, 60, 61, 62, 63] have been explored to produce single photons. Existing time-bin-based QRNG models utilize coherent photon sources in conjunction with attenuators to produce random numbers from exponential [15, 14, 16] or uniform [17] distributions if the underlying optical devices perform ideally and the source generates perfectly single photon. But, in practice, this is not the case. Nie et al. [17] shows that, for a photon arrival-time-based QRNG, the uniformity and hence the min-entropy depends on device parameters such as expected photon number within a fixed time interval as well as on the detection efficiency. Therefore, it is important to explore photon statistics in practical scenarios in more detail.
The QRNG output theoretically follows a perfectly uniform distribution. However, due to the imperfections of the underlying devices, the raw QRNG data may deviate from uniform randomness and therefore require classical post-processing. One may note that deterministic classical post-processing does not increase the randomness of the QRNG output, it only compresses the raw data with an aim to extracting the min-entropy present in it at the cost of the randomness generation rate.
1.1 Motivation
@@
Given that both PRNGs and QRNGs produce sequences of 0’s and 1’s, merely analyzing these sequences is not sufficient to definitively differentiate between the two [64]. Statistical tests like NIST [65], Dieharder [66], AIS-$31$ [67], ENT [68] etc. are available to test the randomness of an RNG. However, from the results of these tests, it is not possible to say whether an RNG is quantum or not. So a pertinent question arises:
@@
Also, there are several schemes of device-independent QRNG (DI-QRNG) [69, 70, 71, 72, 73, 74, 75] and certified randomness [69, 76, 77, 71, 78, 79, 80, 19, 81, 82, 83] based on Bell violation, non-separability, etc. [71, 84, 85, 86]. Though they do not rely on trusted device assumptions and do not require any post-processing, the required resources (that is, entangled qubits) are too costly. More recent works [87, 88] attempt locally certified randomness without entanglement, but they require random seeds and some kind of bootstrapping that may be considered analogous to classical post-processing.
On the other hand, the optical trusted device QRNG [10] relies on the assumption that the devices used for extracting randomness and the source are well-characterized and no one can change any configuration. This means that no third party can access the QRNG or influence its results. In contrast, single-photon-based QRNGs can still be verified using photon-count statistics [52, 53, 89, 90] improving the security. A detailed comparison between trusted QRNG and DI-QRNG is given in Appendix A. In this work, we focus on the QRNG of the optical trusted device type along with a two-fold statistical method to verify its quantumness.
@@
Side-channel attacks [91, 92, 93] can be performed in the post-processing step to gather information about the random number. Therefore, to reduce the compression amount of raw data and to prevent the data from side-channel attacks, the amount of post-processing should be minimal to ensure the maximum entropy in raw data with high speed and reduce the possibility of side-channel attacks on post-processing. To achieve this, the raw output number (before post-processing) generated by QRNG should be as close to uniform random as possible. Nie et al. [17] provided a lower bound for the min-entropy of the raw data in the presence of device errors. Depending on this min-entropy, a suitable post-processing algorithm can be chosen to retain the entropy from the quantum source. As all the practical devices are prone to error, this raises an important question:
Q2: How to increase the min-entropy of the raw data in presence of the device errors?
@@
Q1 and Q2 regarding quantumness and randomness are crucial from the manufacturer’s perspective. Besides these, the random number generation rate and the cost of the QRNG are also important parameters from the user’s perspective. To achieve a comprehensive treatise, we analyze QRNG from both perspectives. Depending on the users’ requirements, the trade-off between the randomness of the quantum source, the random number generation rate, and the cost of the QRNG can vary. For example, in the case of secure communication, the randomness from the quantum source should have the highest priority, and also the generation rate should be high. On the other hand, in token generation or lottery games, a low rate of random number generation would be sufficient. Therefore, to fulfill the user’s requirement, the answers to the following questions are required.
@@ 1.2 Our Contributions
In this paper, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. A similar analysis can also be performed for other QRNG models. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level.
@@
As another contribution, also from the manufacturer’s perspective, we derive the photon statistics for two-photon arrival-time-based QRNGs in practical scenarios. For the practical scenario, we consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
@@
From the user’s perspective, we see that the maximum entropy from the quantum source may not be achieved due to some practical constraints related to the random number generation rate, the timing error in registering the photon detection time and the cost of the QRNG. In particular, we show the following.
@@
  • If the expected number of photons within a unit of time is small, the min-entropy from the quantum source would be high. However, it would increase the cost of the QRNG, and it would also reduce the random number generation rate.
@@
  • Although a small external reference cycle of the QRNG increases the min-entropy from the quantum source, to minimize the timing error, it should be higher than a specific value. This reference cycle does not have much effect on the cost or random number generation rate of the QRNG.
@@
  • A detector with less efficiency increases the min-entropy from the quantum source and also reduces the cost of the QRNG. However, it also reduces the random number generation rate.
@@
Finally, we simulated raw data depending on photon statistics we have derived and performed statistical tests like NIST [65], Dieharder [66], AIS-$31$ [67] and ENT [68] on those raw data. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources. Nie et al. [17] performed NIST tests after post-processing and showed that the post-processed data passed all of the 15 tests from the NIST suite.
@@
This paper is structured as follows. In Section 2 we discuss the quantumness of a single-photon source. In Section 3 and 4 we explore single-photon-based QRNGs and coherent pulse-based QRNGs respectively and mention their weakness in present days. We also discuss noisy beam-splitters and detectors in the case of single-photon-based QRNG in Section 3. Section 5 reveals our two-fold proposal to check the quantumness of a photon-based quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section 6 to answer Q2. In this section, we have also mentioned how our findings are different from the work of Nie et al. [17]. Section 7 answers Q3 by separately analyzing the effects of the parameters mentioned in that question. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section 6 are given in Section 8. Finally, in Section 9 we conclude our work.
2 Quantumness of a Single-photon Source
In this section, we describe what we mean by quantumness in the case of a single-photon source. We first discuss photon statistics and then revisit the existing procedure to distinguish quantum light from classical light using such photon statistics. Here, we validate the quantumness of a photon source by its statistical closeness to an ideal single-photon source. Another way of defining and validating quantumness is through Bell test, quantum discord, non-separability, etc. [71, 84, 85, 86] for QRNGs that are based on entangled photon source. As our point of interest is the single-photon source which does not have any entanglement, we are not concerned with these tests in our current work.
@@ 2 Quantumness of a Single-photon Source
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols [4, 5]. In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, photon-counting statistics play an important role in determining whether or not light falls in the quantum regime.
@@
The distribution of the single-photon-counting statistics can be divided into three classes [53]:
@@
Classical wave theory of light follows Poissonian and super-Poissonian statistics [53]. On the other hand, the sub-Poissonian nature of a photon source cannot be explained by the classical wave theory, making it a clear indication of its quantum nature [89, 90]. A detailed discussion is given in Appendix B.
@@
$P(n)=\begin{cases}\frac{e^{-\alpha\mu}(\alpha\mu)^{\alpha n}}{(\alpha n)!},&\text{if }\alpha n\text{ is integer,}\\ 0,&\text{otherwise}\end{cases}$ (1)
for $n=0,1,2,\dots$ and some positive real numbers $\mu,\alpha$. Then the mean would be $\mu$ and variance would be $\frac{\mu}{\alpha}$. Therefore, this denotes a sub-Poissonian, Poissonian, or super-Poissonian distribution for $\alpha>1,\alpha=1,\alpha<1$ respectively.
Thermal light follows Bose–Einstein distribution [53], given by
@@
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is the creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution [94] given by
@@
Antibunching property can also be used to distinguish quantum light from classical light. Detailed discussion about this is included in Appendix B.
3 QRNG based on Single-photon Source
An ideal QRNG, based on single photons, uses $/50$ 50 beam splitter [95], where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Fig. 1.